OpenStudy (anonymous):

Find the slope of the secant line passing through the points (1, f(1)), (1 + h, f(1 + h)) when f(x) = 2x2 − 2x − 4.

OpenStudy (anonymous):

well the slope would be \[[f(1)-f(1+h)]/h\] or \[[-4-(2[1+h]^{2}-2[1+h]-4)]/h\] which is equal to \[-3-2h\]

OpenStudy (anonymous):

actually I made a mistake first term in the equation is suppose to written as2x^2-2x-4

OpenStudy (anonymous):

i realized that :) its still the right answer

OpenStudy (anonymous):

but the answer choice do not have -3-2h, its either a) slope = 2h + 2 b) slope = 4h − 6 c) slope = 2h − 2 d) slope = 2h + 6 e) slope = 4h − 2 f) slope = 4h + 6

OpenStudy (anonymous):

oh yes.. the answer is a).... apologies I didn't simplify right :P.. \[[-4-(2[1+h]^{2}-2[1+h]-4]/h\] actually is \[-2-2h\] which is still wrong but off by a factor of -1 because I the denominator of the fraction is -h not h... so many mistakes.. sorry about that

OpenStudy (anonymous):

that ok, thanks for help

OpenStudy (anonymous):

Wait, can you step by step show how you calculate -2-2h because I'm still now getting it myself?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

So when finding the slope of something you use the slope formula (y2-y1)/(x2-x1) those are not squared they are subscripts f(1+h)-f(1)/(1+h)-1 the ones will cancel out now those f values in every where you see an x in your polynomial equation

OpenStudy (anonymous):

I had it all written out then I was logged off, sorry i will try to write it out again

OpenStudy (anonymous):

its ok, the same happen to me before, lol

OpenStudy (anonymous):

2+4h+2h^2-2-2h-4-2+2+4/h, now cross out 2h+2h^2/h now factor out 2h 2h(1+h)/h cross out the h's 2(1+h) or 2+2h or 2h+2

OpenStudy (anonymous):

which is a. :)

OpenStudy (anonymous):

did you use the slope formula?

OpenStudy (anonymous):

yes I did, i showed that in my first comment

OpenStudy (anonymous):

It might help if you write out what i did so you can see it on paper

OpenStudy (anonymous):

it just hard to tell because of the way its on the work is on the computer if you were using the slope formula?

OpenStudy (anonymous):

yea its hard to type math formulas but i did use it.